# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v3

@@ Abstract
diversity , ensemble learning , multiple hypotheses prediction , predict-then-optimize , robust portfolio optimization
@@
[…]
[inst1] organization=University of Reading, Department of Computer Science, addressline=Whiteknights House, city=Reading, postcode=RG6 6UR, country=United Kingdom, email=a.j.rodriguezdominguez@pgr.reading.ac.uk
@@ 1 Introduction
The problem of portfolio allocation with a quantitative approach was introduced by Harry Markowitz in 1952 in the so-called Modern Portfolio Theory (MPT) (Markowitz, 1952). It consists of two phases: asset selection and portfolio weight optimization. The asset selection phase decides which assets from a larger set, such as a market, will be included in the portfolio, which has a fixed number of assets but not predetermined choices or weights. The portfolio weight optimization phase determines the percentage of capital allocated to each selected asset. It also has the merit of introducing the concept of diversification as a fundamental component in risk management for investments. In this specific case, diversification is induced by the covariance matrix (Markowitz, 1952). Since then, a large number of methods known as ”plug-in” (or data-driven) approaches have been developed which use optimization inputs with improved estimates of future returns (Best and Grauer, 1991; Broadie, 1993; Giglio et al., 2022; Kelly and Xiu, 2023).
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022; Kelly and Xiu, 2023). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
The PSEM can be defined as a map $\mathcal{E}:\mathcal{X}\to\boldsymbol{r}$, which aggregates predictions as
$\mathcal{E}(\boldsymbol{x})=g\left(f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{1}),f_{\boldsymbol{\theta}_{2}}(\boldsymbol{x}_{2}),\dots,f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{M})\right)$
from multiple individual predictors or hypotheses $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, to produce portfolio predictions $\boldsymbol{\hat{r}}$. Here, $\mathcal{X}$ is the input space, $\boldsymbol{r}$ is the output space, $f_{\boldsymbol{\theta}_{j}}:\mathcal{X}\to\mathcal{Y}$ represents the $j$-th predictor focused on forecasting the time series returns of the $j$-th component of the portfolio, $\boldsymbol{x}_{j}$, and $g:\mathcal{Y}^{M}\to\boldsymbol{r}$ is the aggregation function (e.g., averaging, voting, or weighted combination). Individual predictors are trained with stochastic gradient descent (SGD) with a parameter $\varepsilon\in[0,1]$ that controls their diversity in learning at each step $i$ so that, for the $j$-th predictor:
$\boldsymbol{\theta}_{j}=\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta\left(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})\right)$
where $y_{ij}=x_{(i+\tau)j}$ with time lag $\tau>0$ is the target, $\eta_{j}$ is the learning rate, and the loss function $\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j}),\boldsymbol{y}_{j})$ combines a squared error term and regularization with parameter $\lambda_{p}$. The indicator function $\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})))$, which adjusts the extent to which each predictor’s parameters are updated, is defined as:
$\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})))=\begin{cases}1-\varepsilon,&\text{if }y_{ij}\in\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij}))\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$
Here, $\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the $j$-th predictor’s output is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the top-performing predictors, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble.
After the training, the predictions of all predictors form a structured data set:
$\boldsymbol{\hat{x}}^{T}=\boldsymbol{\hat{x}}(\varepsilon)^{T}=\begin{bmatrix}f_{\boldsymbol{\theta}_{1}}(x_{11})&\cdots&f_{\boldsymbol{\theta}_{M}}(x_{1M})\\ \vdots&\ddots&\vdots\\ f_{\boldsymbol{\theta}_{1}}(x_{N1})&\cdots&f_{\boldsymbol{\theta}_{M}}(x_{NM})\end{bmatrix}$ (3)
@@
$\boldsymbol{\hat{r}}=g(\boldsymbol{\hat{x}}(\varepsilon)^{T})\boldsymbol{w}$ (4)
with ensemble parameters $\boldsymbol{w}$, which, when properly scaled, are equivalent to portfolio weights. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
Figure 1 shows a sketch of the model architecture, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $(\varepsilon)$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the s-RBFN shown in this figure and trained using least squares, described in detail next (Dominguez et al., 2025).
@@
In the next section, the PSEM optimization in the case of an s-RBFN model is described in detail for obtaining the portfolio weights.
@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
In this case, the structured dataset is used as input to a radial basis function network (RBFN), where each $j$-th predictor $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$ is associated with a specific basis function $\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. The basis function $\phi_{j}$ is applied elementwise to the entries of the $j$-th input vector of predictions $\boldsymbol{\hat{x}}_{j}(\varepsilon)\in\mathbb{R}^{N}$, with $\mu_{j}$ and $\sigma_{j}$ representing the center and scale parameters specific to the $j$-th predictor. This defines a transformation map $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon)):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$, where $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon))_{ij}=\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. For example, the Gaussian basis function $\phi_{j}\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{\hat{x}}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ can be used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{\hat{x}}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
$\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{T}\boldsymbol{w}=\left[\begin{matrix}\phi\left(\boldsymbol{\hat{x}}_{11}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\boldsymbol{\hat{x}}_{1M}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \vdots&\ddots&\vdots\\ \phi\left(\boldsymbol{\hat{x}}_{N1}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\boldsymbol{\hat{x}}_{NM}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \end{matrix}\right]\left[\begin{matrix}w_{1}\\ \vdots\\ w_{M}\\ \end{matrix}\right]$ (5)
@@
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain each within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\boldsymbol{w}=1$ and $0\leq\boldsymbol{w}\leq 1$. If (6) is reduced to:
@@
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}$. The target portfolio time-series returns is given by:
$\boldsymbol{r}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (8)
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (6), $\boldsymbol{r}$ must be given by (8). To ensure in the PSEM that the optimization is guided by the structure of the constituents’ hypothesis prediction space and the degree of diversification within the structured dataset of predictions already incorporated into $\boldsymbol{P}$, a canonical basis of hypotheses is selected. For that, the weights in equation (7) are selected such that the target portfolio $\boldsymbol{r}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not on other pre-specified targets:
@@
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from Modern Portfolio Theory (Markowitz, 1952), shown on the left, and a structured Radial Basis Function Network (s-RBFN) optimized via least squares (Dominguez et al., 2025), shown on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}$, and the target portfolio returns, denoted as $\boldsymbol{r}$ in the left case are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, and the s-RBFN model basis functions, $\boldsymbol{\phi}$ and $\boldsymbol{\Phi}$.

### v4

@@ Abstract
diversity , ensemble learning , multiple hypotheses prediction , portfolio optimization , predict-then-optimize , structured models
@@ 1 Introduction
The problem of portfolio allocation with a quantitative approach was introduced by Harry Markowitz in 1952 in the so-called Modern Portfolio Theory (MPT) (Markowitz, 1952). It consists of two phases: asset selection and portfolio weight optimization. The asset selection phase decides which assets from a larger set, such as a market, will be included in the portfolio, which has a fixed number of assets but not predetermined choices or weights. The portfolio weight optimization phase determines the percentage of capital allocated to each selected asset. It also has the merit of introducing the concept of diversification as a fundamental component in risk management for investments. In this specific case, diversification is induced by the covariance matrix (Markowitz, 1952). Since then, a large number of methods known as ”plug-in” (or data-driven) approaches have been developed which use optimization inputs with improved estimates of future returns (Best and Grauer, 1991; Broadie, 1993; Giglio et al., 2022).
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
The PSEM framework is presented as a tool to model a portfolio as a structured ensemble for prediction across multiple hypotheses. This enables a model-centric approach that links predictive generalization within these models to the out-of-sample risk management of a portfolio. Formally, the PSEM can be defined as a map $\mathcal{E}:\boldsymbol{x}\in\mathbb{R}^{N\times M}\to\boldsymbol{r}\in\mathbb{R}^{N}$ which aggregates M predictions from multiple individual predictors or hypotheses {$f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, to produce portfolio predictions for a given asset selection time series return $\boldsymbol{x}$:
$\mathcal{E}(\boldsymbol{x})=g\left(f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{1}),f_{\boldsymbol{\theta}_{2}}(\boldsymbol{x}_{2}),\dots,f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{M})\right)=\boldsymbol{\hat{r}}$ (3)
where $g(\cdot)$ denotes the centroid combiner rule (e.g., averaging, voting, or weighted combination) for the ensemble (Wood et al., 2024). Individual predictors are trained with stochastic gradient descent (SGD) with a parameter $\varepsilon\in[0,1]$ that controls their diversity in learning at each step $i$ so that, for the $j$-th predictor, the parameter update rule follows:
$\boldsymbol{\theta}_{j}=\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta\left(f_{\boldsymbol{\theta}_{j}}(x_{ij})\right)$ (4)
where $y_{ij}=x_{(i+\tau)j}$ with time lag $\tau>0$ is the target, $\eta_{j}$ is the learning rate, and the loss function $\mathcal{L}$ combines a squared error term and regularization with parameter $\lambda_{p}$. The indicator function $\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$, which adjusts the degree to which each predictor’s parameters are updated, is defined as:
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))=\begin{cases}1-\varepsilon&\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})>\mathcal{L}(f_{\boldsymbol{\theta}_{k}}(x_{ik}),y_{ik})\ \ \forall k\neq j\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$ (5)
$\delta(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the output of the $j$-th predictor is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the predictors that perform the best, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble.
After training, the predictions of all predictors form a structured data set:
$\boldsymbol{\hat{x}}(\varepsilon)=\begin{bmatrix}f_{\boldsymbol{\theta}_{1}}(x_{11})&\cdots&f_{\boldsymbol{\theta}_{1}}(x_{N1})\\ \vdots&\ddots&\vdots\\ f_{\boldsymbol{\theta}_{M}}(x_{1M})&\cdots&f_{\boldsymbol{\theta}_{M}}(x_{NM})\end{bmatrix}$ (6)
@@
$\boldsymbol{\hat{r}}=g(\boldsymbol{\hat{x}}(\varepsilon)^{T})=\boldsymbol{\hat{x}}(\varepsilon)^{T}\boldsymbol{w}$ (7)
with ensemble parameters $\boldsymbol{w}$, which, when properly scaled, are equivalent to portfolio weights. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in MPT (Markowitz, 1952).
Figure 1 shows a representation of the PSEM optimization, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $\varepsilon$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the Structured Radial Basis Function Network (s-RBFN) shown in this figure and trained using least squares (Dominguez et al., 2025).
@@
In the next section, the PSEM optimization using least squares is detailed for the case of the s-RBFN model. Additionally, the equal-weighted portfolio is presented as the target return in the PSEM optimization for all cases, serving as a canonical basis of hypotheses.
@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
In this case, the structured dataset is used as input to a radial basis function network, where each $j$-th predictor $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$ is associated with a specific basis function $\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. The basis function $\phi_{j}$ is applied elementwise to the entries of the $j$-th input vector of predictions $\boldsymbol{\hat{x}}_{j}(\varepsilon)\in\mathbb{R}^{N}$, with $\mu_{j}$ and $\sigma_{j}$ representing the center and scale parameters specific to the $j$-th predictor. This defines a transformation map $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon)):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$, where $\boldsymbol{\Phi}(\boldsymbol{\hat{x}}(\varepsilon))_{ij}=\phi_{j}(\hat{x}_{ij}(\varepsilon),\mu_{j},\sigma_{j})$. For example, the Gaussian basis function $\phi_{j}\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{\hat{x}}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ can be used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{\hat{x}}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
$\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{T}\boldsymbol{w}=\left[\begin{matrix}\phi\left(\hat{x}_{11}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\hat{x}_{1M}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \vdots&\ddots&\vdots\\ \phi\left(\hat{x}_{N1}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left(\hat{x}_{NM}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \end{matrix}\right]\left[\begin{matrix}w_{1}\\ \vdots\\ w_{M}\\ \end{matrix}\right]$ (8)
@@
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, ensuring they sum to one and remain each within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\boldsymbol{w}=1$ and $0\leq\boldsymbol{w}\leq 1$. After the optimization in (9), the target portfolio time-series returns would be given by:
$\boldsymbol{r}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (10)
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (9) $\boldsymbol{r}$ must be given by (10). To ensure in the PSEM that optimization is guided by the structure of the hypothesis prediction space of the constituents and the degree of diversification within the structured dataset of predictions already incorporated into $\boldsymbol{P}$, a canonical basis of the hypotheses is selected. where $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}$ consequently reducing (9):
@@
The weights in equation (11) are selected such that the target portfolio $\boldsymbol{r}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not on other pre-specified targets:
@@
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from MPT (Markowitz, 1952), shown on the left, and the s-RBFN optimized via least squares (Dominguez et al., 2025), shown on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Figure 2: Systems representation comparison of the Mean-variance framework to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the PSEM framework with the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, $\boldsymbol{x}$, and the target portfolio returns, $\boldsymbol{r}$, in the left case are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, and the s-RBFN model basis functions in matrix form $\boldsymbol{\Phi}$.
@@
The next section describes the methodology used to conduct portfolio allocation experiments, encompassing both the asset selection and weight optimization stages, from the perspective of multi-hypothesis prediction with structured models. This approach seeks to uncover connections between diversity in asset selection and out-of-sample portfolio diversification, identify new limits on portfolio diversification introduced by this additional source of variability, and evaluate the ability to control these effects parametrically prior to weight optimization.